Suppose that 5,700 is invested at 5% annual intrest rate, compounded monthly how much money will be in the account in 4 months? and in 3 years [round to the nearest cent]





how many years will it take 7,000 to grow to 11,500 if it is invested at 9% compounded quarterly? and compounded daily?

Answer by rfer(16322)

A=5700(1+0.05/12)^4
A=5700(1+0.05/12)^12*3
11500=7000(1+0.09/4)^4n
11500=7000(1+0.09/365)^365n